OpenStudy (anonymous):

1. Which of the following is the vertex of the quadratic equation y = –4(x + 6)^2 + 2?

OpenStudy (cwrw238):

for y = c(x + a)^2 + b vertex is ( -a, b)

OpenStudy (anonymous):

so what do i do basically do i plug it in ? if so where

OpenStudy (cwrw238):

compare the two equations and replace a and b y = c(x + a)^2 + b y =-4(x + 6)^2 + 2

OpenStudy (anonymous):

okay ive got it now thank you .

OpenStudy (cwrw238):

yw